OpenStudy (anonymous):

Find the number of units x that produces a maximum revenue R. R = 114x^2 − 0.04x^3 x = _______units

OpenStudy (anonymous):

So far I have taken the 1st derivative and 2nd derivative I have R'=228x-.12x^2 R" =228-.24x Now I set R' to equal 0 which is R' 228x-.12^2=0 I'm not sure how to get the critical point?

OpenStudy (anonymous):

Somebody PLEASE help me??/

OpenStudy (amilapsn):

zeros of R' are the critical points...

OpenStudy (anonymous):

I dunno how to get the number for x.... it must be a large number because i can't figure out x...

OpenStudy (amilapsn):

Factorize R', then it's easy to figure out zeros...

OpenStudy (anonymous):

R' x(228-.12x)??? I don't see what x can be still

OpenStudy (imstuck):

What you have after you find your first derivative is 228x-.12x^2.

OpenStudy (amilapsn):

What are the x values in the critical points?

OpenStudy (imstuck):

in other words, you have this:\[R'=-.12x ^{2}+228x\]I would multiply by 100 to get rid of the decimal, like this:

OpenStudy (amilapsn):

They are the x values when R' = 0

OpenStudy (imstuck):

\[R'=-12x ^{2}+22800x\]Of course I would also switch the signs:\[R'=12x ^{2}-22800x\]Now factor out an x and get this:\[R'=x(12x-22800)\]x = 0 or 12x-22800=0. 12x=22800 and x = 1900

OpenStudy (imstuck):

x=1900 units.

OpenStudy (anonymous):

ok, that's what i was looking for IMSTUCK.. I understand how to do the whole problem except the whole - decimal number kept throwing me off. Thank you for the steps again though. I will keep that in mind when i have - decimal numbers

OpenStudy (anonymous):

all i have to do to check is plug 1900 in for x and i should get 0 :)